### v5

# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative LLM-as-a-Judge

Abstract
LLM-as-a-Judge offers a promising alternative to human judges across various tasks, yet inherent biases, particularly position bias—a consistent preference for answers based on their position in the prompt—compromise its effectiveness. Our study proposes a systematic framework to investigate the position bias of LLM judges in the pairwise comparative scenario in terms of repetition stability, position consistency, and preference fairness. In pairwise comparative settings, LLM judges are prompted to choose the better candidate’s answer to the task question, where the task question and a pair of candidate answers are provided. Then the prompt with swapped order of candidate answers is given to the same judge again. If the judge prefers the same winner, it is denoted position consistent; otherwise, if different winners are selected after swapped settings, position bias is observed.
To comprehensively understand the position bias of LLM judges, our experiments cover 12 judges across MTBench and DevBench with 22 tasks and around 40 answer-generating models, resulting in over 100,000 evaluation instances. Further exploration of the Judge-level, Candidate-level, and Task-level impactful factors on position bias enhances the understanding of this bias. Through comprehensive investigations, we validate that the position bias of capable LLM judges is not merely random variations. We also observe significant variance of position bias across judges and tasks, highlighting trade-offs of positional consistency, preference fairness, and practicality when selecting task-specific judges. Among impactful factors, position bias is largely influenced by the quality gap of given candidate answer pairs for evaluation. However, the lengths of the prompt components, including input questions, candidate answers, and the complete prompt, have minimal impact on position bias. In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs. These insights guide the selection of optimal judge models, enhance benchmark design, and lay the foundation for future research into effective debiasing strategies, ultimately improving the reliability of LLM evaluators for more robust, unbiased evaluation systems.
Provocations
Despite advancements in LLM-as-a-Judge and prevalent attention on position bias, the understanding of such bias is still limited and requires a more systematic evaluation. Enhancing the understanding of the impactful factors on position bias could inspire more effective mitigation strategies, laying the foundation for more reliable, robust, and unbiased evaluation systems. To this end, our study investigates aspects that are underexplored by prior works to offer a systematic evaluation framework for the position bias of LLM judges.
Repetition Stability
Validating LLM judges are repetition stable is a primary challenge that is often ignored by previous studies. They trusted the one-time generated judgments of the LLMs and conducted further analysis, implicitly assuming that the judgments represent the judge’s capabilities. However, the judgment choices and consequently observed position bias may be due to random variations. In other words, without verifying that LLM judges are stable in their choices across multiple trials, it is uncertain whether position bias stems from positional shift or is merely randomness. To address this gap, our study utilizes repetition stability to validate capable LLMs before further investigations on position bias.
Preference Fairness
While position bias is prevalent across most tasks and widely studied by literature, the main focus is on position consistency, as it is the most straightforward measurement of position bias. However, the preference directions when position bias occurs are largely underexplored by prior works. In this study, we show that previous measurements are either sensitive to the dataset size or overly penalizing highly consistent judges. To address this problem, we propose a more granular and scalable measurement that combines the strengths of both methods to assess preference fairness. We further stress the trade-off between position consistency, preference fairness, and practicality when choosing the task-specif judges, as position bias largely varies by task and judge.
Answer Quality Gap
From intuition, the difficulty of judging a pair of candidate answers (i.e., choosing the winner that better answers the task question) largely depends on the disparities of the answer pairs in quality. However, we are the first study to quantify the answer quality gap and investigate its correlation with position bias. Our investigations verify that as the answer quality gap enlarges, implying one candidate’s answer is significantly better than the other, the judges become more constant and fair. The findings not only benefit the interpretation of other LLM judge biases like verbosity/length bias and self-enhancement bias, but also benefit future benchmark designs for more social goods.
To summarize, this study proposes new metrics and studies more impacting factors on position bias that are underexplored by prior studies, offering a more comprehensive and systematic framework for understanding and evaluating the position bias of LLM judges. We stress the need to have a deeper understanding of the complex bias before moving to solutions. More detailed analysis, insights, findings, and contributions can be found in the Appendix.
[…]
[appendices]
Appendix Table of Contents
[…]
[appendices]0
Appendix
Appendix A Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global attention and stimulating substantial research into their applications. Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations [1, 2]. To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments[3, 4, 5], inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1. This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations [6], for any score-based and relation-based settings [7]. Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias. However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks [8, 9]. Moreover, the limited understanding of the bias itself has exacerbated the problem.
Prior studies on position bias implicitly assume that this bias is not due to random variations. In other words, these studies assume that LLM judges produce consistent results across repeated trials. However, without validating this assumption, evaluations may become less reliable. As shown in Fig. 1, the occurrence of position bias can be observed by chance even for the same set of tasks and candidate models. Thus, it is unclear whether the observed position bias actually arises from the position of the prompt components or is merely due to random chance. Therefore, when exploring the position bias of LLM judges, it is necessary for the judges to be stable across repetitions to yield nontrivial results.
Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge. When the positions of model responses are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same response. We observe repetition instability if the same LLM judge gives different judgments across repetitions. Only when an LLM judge has high repetition stability can we validate that its position bias is not due to random variations.
Another key aspect of position bias, preference fairness, has also been under-explored by previous works. Most studies have primarily focused on positional consistency to measure how frequently position bias occurs across a set of evaluation instances. However, the direction of biased preference (i.e., primacy or recency) is also insightful. Fig. 2 illustrates cases where a pair of judgments with swapped prompts become positionally consistent, inconsistent, primacy-preferred, or recency-preferred. We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations. Specifically, judgments are considered to have a fair preference if they are positionally consistent or if the preference directions are evenly distributed. In contrast, an LLM judge is considered to have an unfair preference if it consistently favors responses that appear either first or last in the prompt. Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
Figure 2: Demonstration of positionally consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better response. A judgment is consistent if it selects the same winner across different prompt positions. Primacy/recency indicates a preference for the first/last model response, respectively. Positionally inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment. We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring [3, 10], making it a practically preferable choice for investigations.
Our contribution can be summarized as following: (1) introducing a systematic framework for evaluating LLM-as-a-Judge’s position bias in terms of Repetition Stability, Position Consistency, and Preference Fairness; (2) identifying Judge-level, Candidate-level, and Task-level factors that affect position bias; and (3) providing insights into future benchmark design and paving the way for effective mitigation strategies for position bias.
Appendix B Methods & Defintions
LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments. Our study focuses on relation-based assessment, specifically under the pairwise comparison setting, for several reasons. First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields [6]. Second, as noted by [11], LLM and human evaluations are more aligned under the context of pairwise comparison than any of the score-based assessments. Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of positional consistency[3, 10]. Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, by some advanced ranking algorithms [6, 11]. Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
In the context of pairwise comparative assessments, LLM-judges are prompted to choose the better response that answers the question. To accommodate the possibility of a tie, various option modes are introduced. The Two-Option mode restricts judges to choosing the better response between two options, labeled {A} for the first and {B} for the second as shown in Fig. 1. The Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither option is preferable. The Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie. In our experiments, we employ the Three-Option mode for MTBench [3] and the Two-Option mode for DevBench [4], aligning with their respective original frameworks. These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM-judges.
To setup our experiment, as shown in Fig. 1, the system prompt, option mode, task question, answers/responses from the two candidate models (original prompt) are presented to the LLM-judges to select the winning answer. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM-judges, and the candidate models are unaware that their responses will be compared to another model when answering the question. Then, the prompt with answers in a swapped position (swapped prompt) is given to the same judge again, which results in a choice pair. As shown in Fig. 2, if the choice pair consistently favors the same answer regardless of the swapped position, it is considered positionally consistent with a fair preference. Conversely, if the LLM-judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
To systematically investigate position bias in pairwise comparative judging scenarios, we evaluate LLM judges using three key metrics: Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$). While $PC$ has been widely adopted in prior research, we introduce $RS$ to validate the reliability of LLM judges and $PF$ to provide a more comprehensive assessment. Further analysis examines the impact of judge-level, candidate-level, and task-level factors on position bias. Therefore, for each judge, candidate, and task, we calculate $RS$, $PC$, and $PF$ to thoroughly explore the variance in position bias and its impacting factors.
B.1 Evaluation Metrics
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the choices of LLM judges, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,D\}$
where $S$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries. The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
Position Consistency ($PC$) quantifies how frequently a judge model selects the same response after the order of responses is swapped. It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the model selects the same winner. Formally, it is calculated as
$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}}$
where $(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})$ denotes the judgment pair for the $j^{th}$ query, $n$ represents the number of prompt pairs, and $V$ is the set of choice pairs that correspond to position consistency. An example of such pairs of choices can be found in Fig. 2. This formula provides a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain response positions. In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second position" used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 2.
Positionally consistent judgments are fair in preference. The measurements become more complex for positionally inconsistent cases. Previous studies proposed two common ways to measure the preference fairness. A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs [3, 5]. However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like [12, 10] treat positionally inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of positionally inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach is problematic because it overlooks the fact that “positionally consistent judgments are also positionally fair," which can lead to overly penalizing highly consistent LLM-judges. For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred. However, this outcome should be distinguished from cases where all one hundred pairs are inconsistent and biased toward one position, which would also result in 100% preference unfairness.
To overcome these limitations, we introduce a more granular and scalable measurement that combines the strengths of both methods, to assess preference fairness. The $PF$ score is formally calculated by
$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the positionally consistent judgments. Another advantage of this approach is that it quantifies positional preference with a single score, rather than the two separate scores used in previous studies. The $PF$ score is interpreted as follows:
$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Positionally Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
To summarize, our proposed preference fairness score is superior to prior measurements due to its ability to provide a single, comprehensive score that accounts for all evaluation instances, while maintaining its sensitivity across datasets, even when the number of questions and instances varies for each task.
B.2 Factors affecting position bias
Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose responses to the task-specific questions are evaluated by the judges. An asterisk (*) highlights the factors with significant impact on position bias that we found through experiments. ‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s answers to the question, making it both a Candidate-level and Task-level factor. ‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
Factor Judge-level Candidate-level Task-level
Familial Property $\checkmark$* $\times$ $\times$
Answer Quality Gap $\times$ $\checkmark$* $\times$
Task Input Length $\times$ $\times$ $\checkmark$
Task Output Length $\times$ $\checkmark$ $\checkmark$
Prompt Length $\times$ $\checkmark$ $\checkmark$
To investigate the factors influencing position bias in LLM judges, we categorized these factors into three groups: Judge-level, Candidate-level, and Task-level factors. Each group includes specific factors, that we hypothesize, may impact position bias, which we explore through a series of experiments. Table 1 lists the five factors analyzed in this study. By design, our framework allows for the integration and assessment of additional influencing factors.
Among the influencing factors we identified, we chose “familial property” instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments. The model family grouping reflects similar model sizes and training specifics of the model members. The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version. However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable" models sharing similar judging capabilities. More details and discussions about the familial property can be found in Appendix Sec. H, where we examined the LLM mutual agreements to study the similarities and differences across families.
Answer quality gap, based on our findings, is a highly impactful yet under-explored factor in prior research. We define the quality of a model’s answer by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between two candidate model answers to the same question and is considered a Candidate-level (i.e., response provider) factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs, it would prefer the higher-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
Depending on option mode, we regard a tie judgment as‘half-win’ for both candidates when calculating its win rate over another.
Then, the win rates between two candidates on a set of task questions can reflect the answer quality gap between them on the task.
However, if position bias is exhibited, the winner may be inconsistent when the order of candidate responses is swapped in the query. Therefore, we observe the LLM judgments where the absolute winner is consistently chosen in both original and swapped queries (termed “consistent wins”), both responses are deemed equal (termed “consistent ties”), and different winners are selected after candidate responses are swapped in the query. We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent choice pairs ($C_{I}$), respectively. Following [3], we count inconsistent choice pairs as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
Inspired by [3], we define the overall win rate ($owr$) of a model’s response over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties. In contrary to using consistent win rate by [3] (calculated as $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of positionally consistent judgment pairs) to quantify $\delta_{q}$, the adoption of overall win rate incorporates all data points and captures the “comparable quality” scenario, where responses in similar quality might lead to positionally biased judgments, a scenario that the consistent win rate might overlook. More discussions can be found in Appendix. Sec. I.
Appendix C Experiments
C.1 Experiment Settings
In this study, we evaluated twelve up-to-date commercial models from the GPT [13], Claude [14], and Gemini [15] series using our systematic framework. Three Llama models [16], with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and well-known capability in reasoning tasks. However, due to their limited performance on our tasks, as shown in Table 2, we focus our investigation solely on more powerful closed-source models. These models are more widely used and reliable, though we still include open-source models to gain more comprehensive insights.
We adopted the modified MTBench [3] and DevBench [4] datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks. We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired these candidate models with the baseline candidate for evaluation by the LLM judges. The prompt templates we used are identical to those used in the benchmarks. More details about the models, tasks, and prompts can be found in Appendix. Sec. L.
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials. This resulted in 576 and 432 instances per judge for MTBench and DevBench, respectively. The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, approximately 100,000 evaluation instances were analyzed in this study.
C.2 Empirical Results and Interpretations
The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations.
This further necessitates the need for investigating and analyzing preference fairness simultaneously with position consistency. Fig. 3 and Fig. 7 demonstrate that position bias varies by judges and tasks significantly. Together with Table 2, it stresses the trade-off between $PC$ and $PF$ when selecting task-specific LLM judges. Fig. 4 further investigates the impact of answer quality gap and lengths of prompt components on position bias. These analyses are conducted by considering all judges together on MTBench. The detailed analyses for each judge and on DevBench can be accessed from Appendix.Sec.J, and Sec.K.
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Positional Fairness ($PF$). Top 5 performances are marked in bold.“Error" is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format. High error rates and low $RS$ are marked red, implying further evaluations (i.e., $RS$, $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data. Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances. In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations.
Judge MTBench DevBench
$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error
Caude-3.5-sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00
Claude-3-opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00
Claude-3-sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00
Claude-3-haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00
Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13
Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00
GPT-4-turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00
Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
Figure 3: Baseline comparisons of judges across tasks on MTBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be gpt-4-0613 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
Figure 4: Impact of answer quality gap and lengths on position bias. All figures are based on the integration results of all judge models on MTBench. Figures (a) and (b) investigate $PC$ and $PF$ vs. overall win rate, hence the answer quality gap, respectively. Generally, judgments become more consistent and fair as the quality gap decreases. Figure (c) demonstrates the relationship between $PC$ and $PF$, where higher consistency is likely to indicate better fairness. Figure (d)(e)(f) plots the relationship between lengths of prompt components and $PC$. The irregular and inconclusive patterns imply that position bias is not significantly impacted by the length of prompts and their components.
Appendix D Main Findings
Position Bias of Capable Judges are not Mere Random Variations
As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal “Error" rates, generally exhibit $RS$ of over 0.85. The up-to-date and most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, Gpt-4, and GPT-4o, all achieve near-perfect $RS$ of over 0.95 on both benchmarks. These results validate that LLM judgments and the consequently observed position bias are not merely random variations. This increases the confidence that one-time generated judgments by these validated LLMs adequately represent their judging capabilities. Only with this prerequisite, can we ensure further investigations are reasonable and meaningful.
Position Bias Varies by Judge & Task
As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig. 3 for $PC$ comparisons. However, certain models achieve comparable or superior performances than GPT-4 on certain tasks. For example, according to Fig. 3, for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges. Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
Similarly, significant variances of $PF$ across tasks are observed. The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others. Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, roleplay, and writing tasks. Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
Moreover, high position consistency does not guarantee fairness, weighing the trade-offs between position consistency and preference fairness. For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively. In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency. In this sense, sacrificing consistency for more fairness or vice versa can all be practically useful depending on specific needs. This trade-off between consistency and fairness is particularly crucial when the most consistent model is not the fairest.
Therefore, to sum up, the position bias of LLM judges is task-dependent and there is usually a trade-off between position consistency and preference fairness. In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance. The same discoveries can be concluded from DevBench results, as detailed in Appendix.Sec. K.
Position Bias Correlates to Answer Quality Gap
Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section B.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$). Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1. Fig.4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.J.
Similarly, as shown in Fig. 1(b), judgments generally become more positionally fair as $\delta_{q}$ increases. However, the extent is not as significant as for $PC$. Also, the relationship varies by judge, where some LLMs maintain preference fairness regardless of $\delta_{q}$. For example, as shown in Fig.6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
These observations align with the right-arrow shape as demonstrated in Fig.4(c), where there is a general trend that judgments become preference fairer as position consistency increases. It also justifies the reasonability of our quantification of positional fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves. However, exceptions are common, as shown by the individual scatter points of these figures. This indicates that though the answer quality gap significantly influences the position bias of LLMs, other factors also play important roles. Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
Our investigation also helps to explain the observed verbosity/length bias [11, 17, 3] and self-enhancement bias [12, 3] of LLM judges. From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities. Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the answers generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4). In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
Position Bias is Length-Independent
We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the response length of answer-generating models (task output length), and the length of the entire prompt (prompt length). Since the tokenization methods vary by model, we apply the length of the prompt component strings for uniformity and scalability. The analysis from Fig. 4(d)(e)(f) shows that there is no significant and definite relationship between the lengths of prompt components and position bias. Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
Practical Implications for Benchmark Design
We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appedix.Sec. H to gain more insights. Particularly, the disagreement analysis helps to identify the instances where it is easy/difficult for all judges to reach a consensus, reflecting the difficulty of judging them. Based on discoveries of the answer quality gap and LLM agreement/disagreement, our study offers practical implications for evaluator benchmark designs in controlling the difficulty of judging tasks. From our findings, the hard-to-judge instances are the ones where (1) LLMs mostly disagree with each other (2) comparable win rates and minimal quality gaps are observed for the two candidate models (3) prominent position bias exhibits for a majority of judges.
Here is a qualitative analysis of a typical difficult-to-judge example. This instance (MTBench question 120) involved a math problem where both models, vicuna-13b-v1.3 and wizardlm-30b, provided incorrect answers. The problem statement is as follows: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find x such that $f(x)=0$. Methodologically, vicuna-13b-v1.3 erroneously applied the quadratic formula to solve the cubic equation, whereas wizardlm-30b employed an enumeration method but failed to identify the correct answer. This resulted in a 3{A}-4{B}-5{C} choice pattern among the twelve judges, a typical case of significant disagreement.
Qualitatively, We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer. The remaining ones showed preferences either for correctness (which then chose {C} as a tie) or methodological soundness (which then chose {B}, i.e., preferring wizardlm-30b). Notably, in the swapped order setting of the same problem, 8 of the 12 judges encountered position bias. This is a typical illustration of the two responses having a minimal quality gap, leading to difficulty in judging, the likelihood of position bias, and prominent subjectivity during evaluations.
The takeaway message from examples like this is that the number of hard-to-evaluate and highly subjective instances could be reduced for better benchmark designs. Similarly, the trivial-to-judge instances where all judges easily reach a consensus could be reduced among the benchmark population. One could also specify more detailed and definite evaluation criteria to erase the ambiguity of some hard-to-evaluate instances.
Appendix E Conclusion
In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of pairwise comparative LLM-as-a-Judge in terms of repetition stability, position consistency, and preference fairness. We focus on position bias for its prevalent occurrences and significant impact on the reliability of LLM judges’ evaluations. From prior works, the pairwise comparative assessment offers the most consistent and reliable judgments compared to other types of LLM-as-a-Judge and hence becomes our concentration for practicality. The study starts by investigating the repetition stability of LLM judges to validate that the judgments and consequently observed position bias are not mere random variations. Then, in addition to the traditional measurement of positional consistency, we propose a systematic way to quantify the preference fairness of LLM judges. Our comprehensive evaluation of twelve judge models on two benchmarks across 22 tasks reveals significant variance of position bias by judge and task. This stresses the trade-offs between position consistency, preference fairness, and cost-effective practicality when choosing the optimal judges for specific tasks.
The study on the Judge-level, Candidate-level, and Task-level impactful factors on position bias further enhances the understanding of this bias. For example, we are the first the make a direct connection between answer quality gap and position bias, where our discovery reveals that the quality gap between candidate responses to the task question significantly influences the positional consistency and preference fairness of LLM judges. However, position bias is length-independent, meaning that the lengths of prompt components have minimal impact on position bias. In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs. To summarize, insights from this study enhance the understanding of position bias and lay the groundwork for more robust, unbiased evaluation systems.
Appendix F Contribution & Limitation
Contributions
Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge. Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through positional consistency and positional fairness. We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components. Particularly, our study is the first to establish a direct connection between quantified answer quality disparities and position bias, highlighting the urgent need for a comprehensive understanding of how position bias affects LLM judges. Additionally, the mutual agreement and disagreement analysis of LLM judges are insightful for identifying the difficulty of evaluation instances and beneficial for benchmark assessment and designs. All findings from our study enhance the understanding of the position bias of LLM-as-a-Judge, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Social Benefits Our comprehensive understanding of position bias in LLM-as-a-Judge provides multiple benefits to the community. First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation. Our focus and investigation on positional fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability. Moreover, our emphasized trade-offs between positional consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work. Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results. Besides, the LLM agreement/disagreement analysis in Appendix.5 provides an applicable tool to identify the difficulty of judging certain instances. Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias. Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare [18], instruction following and prompting [1, 9], multimodal assessment [19], and recommender systems [20].
Strength
Our framework of systematically understanding position bias has the following strengths: (1) scalability and ease of use - our proposed framework is scalable to various judges, models, tasks, and judging types. It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works. (2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, positional consistency, and positional fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias. (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
Limitations
Due to computational limits, our study can be potentially extended to provide more comprehensive insights. (1) We only study the pairwise comparative assessment of LLM-as-a-Judge for its best consistency, ease of use, and scalability, but more types of judging can be explored, such as pairwise scoring and listwise ranking. (2) We only study 12 up-to-date commercial LLM judges for systematic investigations. Llama 7b-13b models that we choose as representatives of open-source models of similar sizes, face extreme difficulty in judging the instances explored in this study. Due to limited resources and considering practicality, we did not experiment with open-source models with larger sizes. However, further exploration may include analysis of these state-of-the-art open-source and fine-tuned judge models using our proposed framework. (3) We only study position bias on two benchmarks, MTBench and DevBench, which is still limited despite 22 tasks and 40 answer-generating models. Also, the number of questions for each task is limited. Further studies can be extended to more benchmarks to have a broader understanding of position bias across benchmarks, tasks, and questions. (4) Due to data accessibility, we are unable to investigate Judge-level factors in terms of model parameter sizes and architectures, which we expect will have a significant impact on position bias. Alternatively, we consider the family property to account for Judge-level factors by grouping models with similar properties (e.g., approximate parameter size) together. We also investigate variances of position bias across individual judges. However, future exploration on these model properties, especially those available on open-source models, will provide more insights.
Future Work
Besides addressing the limitations, the following future work, built on findings from our study, could provide more insights: (1)prompt setting - in this study, we apply the exact default prompt settings of MTBench and DevBench. However, the positional order and prompt style of not just the model-generated responses but also system prompt components (e.g., agent role assignment, mode of judging, direct mention of not making biases) may also have an impact on the extent to which LLM judges make position bias. (2) open-source insights: Further investigations of open-source and fine-tuned models on easier-to-evaluate benchmark may provide more insights into the training stages or model structures that are affecting or causing the position bias. Then our framework offers better control on the other impactful factors such as the answer quality gap to make these further investigations more accurate.
Appendix G Related Work
G.1 LLM-as-a-Judge
In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications. For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature [1]. However, it lacks scalability and reproducibility [2].
As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks [3, 4, 21, 10, 22, 23, 24, 25, 26, 11, 1, 27, 28, 29, 30, 2, 31, 17, 32, 33, 34, 35, 36, 37, 38, 7, 12, 39], including open-ended story generation [30], adversarial attacks [21], summarization [2], machine translation [29], and instruction following [1].
These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles [3, 40, 41, 21, 28, 10, 6, 22, 42, 5, 43, 44, 45, 11, 1, 8, 17, 39, 7, 46, 12, 18, 9, 19]. As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation. Therefore, cost-effective LLM judges are needed to assess these responses efficiently. LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
In many of these tasks, LLM judges have shown a high level of agreement with human evaluators [3, 4, 5], yet in some tasks, they are less effective, largely due to inherent biases [28, 21, 40]. Even in cases where agreement is high, judgments may still suffer from biases.
When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method. From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment [6]. By comparative method, it can be either score-based or relation-based [7].
For example, pointwise scoring [17] lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics. Pairwise/Listwise Scoring [2, 5, 19, 3] prompts the LLM judges to score a pair/list of model-generated answers. Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst). Pairwise comparative assessment [2, 19, 3], on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
Except for pointwise evaluation, all forms of LLM-as-a-Judge suffer—or are susceptible to suffering—from position bias due to the intrinsic nature of position and comparison within the prompt structure.
G.2 Position Bias
Previous studies have discovered and investigated multiple types of biases, such as position bias [3, 4, 19, 22, 35, 21, 8, 5, 34, 40, 36, 1, 47, 6, 45, 46, 10, 11, 17, 2, 20, 25, 41, 28, 30, 29, 44, 12, 7, 43], verbosity/length bias [3, 11, 17], self-enhancement bias [12, 3], selection bias [40, 41], and contextual bias [11, 48].
Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source [39, 18], proprietary commercial ones [13, 14, 15], and fine-tuned models [42, 47, 5, 39, 17].
To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content. For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs. Some other studies also use the term position bias [49, 50, 51, 52], but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
Position bias is arguably the most prevalent and impactful bias among all. [9] notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought [53] performance over self-training controls for all biases except position bias. Furthermore, Khan et al. [8] point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
Moreover, there is ongoing debate over whether selection bias originates from position bias. [41] argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.[43] contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
All of these findings underscore the critical importance of addressing the issue of position bias when employing LLM-as-a-Judge.
G.3 Deal with Position Bias
Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts [3, 43, 41, 54, 55]. Also, LLM judges are vulnerable to attacks [43, 44, 46].
Parse data with position bias There are many ways to deal with the position bias. The naive way is to exclude the inconsistent judgments(i.e. if the LLM judge gives a positionally biased judgment on a pair or list of model responses) [3, 44, 45, 39]. While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem. Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging [3, 43, 39, 45] or regard inconsistency as a half-win or tie for relation-based judging [3, 12] after swapping the order of model responses in the prompt. For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B. The latter way of swapping + tie is proposed because intuitively the position bias is more likely to occur when the model responses get evaluated share a similar quality in terms of the evaluating metric. Our quantitative study also verifies this intuition, evidenced by the fact the LLM judge’s positional consistency (the percentage of positionally consistent pairs of judgment) is positively proportional to the answer quality gap between the model responses.
To save the expense of running the experiments more than once using swapped order, in practice, many of these studies also support a random-shuffle option in their code settings such that the baseline model for comparison does not remain in a fixed position.
Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping [20], split-and-merge [7], and multi-agent discussion [12, 8]. However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
Furthermore,[11] suggests that existing calibration techniques designed to reduce bias, including context decomposition [7], order permutation [45, 40], ensembling [56], and batch calibration [48], are insufficient to align LLM evaluators, even with supervised data.
Thus, position bias is a pervasive, significantly impactful, and challenging problem to solve.
Understand before solving We observe that the existing methods maybe ineffective or not enough satisfactory because there is a lack of understanding of the position bias. Although a variety of studies have researched this type of bias in LLM-as-a-Judge, a comprehensive understanding of what factors affect the position bias remains underexplored. In other words, without clarity on the key factors and their quantitative impact on position bias, the efficacy of current and future methods remains uncertain.
For instance, [7] proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%. However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful. In other words, if the quality between the two model responses differs considerably, the LLM judges will make little position bias during judgment and hence is easy to calibrate. This example illustrates how lacking a comprehensive understanding of the factors affecting position bias can lead to overestimating or misjudging the effectiveness of methods proposed to resolve it and improve the performance of LLM-as-a-Judge.
Missing analysis for position bias Besides, the positional preference direction aspect of position bias remains underexplored. Positional preference refers to the specific positions an LLM judge favors when position bias is evident. For instance, if an LLM judge consistently favors the responses that appear first in the prompt, it exhibits a preference for the first position. Our study denotes primacy-preferred and recency-preferred to describe biases toward the first and second positions, respectively. We integrate the measurement by formally defining positional fairness score to quantify the extent of a judge model’s preference towards certain positions.
While prior research includes some positional preference results [3, 45, 5, 12, 10], none have conducted a detailed analysis. They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
Therefore, we propose that mitigating the position bias of LLM-as-a-Judge requires a simultaneous improvement in both positional consistency and positional fairness. A positionally consistent (high positional consistency) and positionally fair (nearly equal preference on primacy and recency when position bias occur) LLM judge is undoubtedly preferable. However, a trade-off between consistency and fairness/preference often occurs in practice. Previous studies that have provided results on positional preference have all shown that even state-of-the-art (SOTA) LLM judges struggle with both positional inconsistency and clear positional preferences.
Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions. MLLM-as-a-Judge [19] conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments. In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness. If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
G.4 Summary of Prior Work
To summarize, LLM-as-a-Judge has a large potential in alternating human judges on a wide range of tasks, in particular subjective evaluation, owing to its cost-effectiveness, high agreement with human judgments, reproducibility, and scalability. However, it suffers from various biases, most notably position bias, which is prevalent across different evaluation tasks, models, and judgment types. Not only is it a type of bias that significantly hinders the improvement and promotion of applying LLM-as-a-Judge, but it is also difficult to solve due to its complexity and lack of understanding in the community.
Existing solutions proposed to address this issue may be ineffective or uncertain in efficacy due to the unclear impact and nature of influencing factors. Additionally, whether the position bias is an essential bias or mere random variation has not been validated from prior works. The positional preference side, reflecting the positional fairness of the LLM judge, also requires comprehensive analysis. In a nutshell, a comprehensive understanding of the position bias of LLM-as-a-Judge is crucial to validate existing and future approaches to address this important problem.
Appendix H LLM Agreement Analysis
Besides the exploration of position bias with a broad lens by average $PC$ and $PF$, instance-wise agreement between LLM judges is also insightful. Even two judges with the same $PC$ and $PF$ scores may not reach consensus on each instance. Therefore, this session investigates (1) what percentage of a set of evaluations do two LLM judges agree on each other? (2) how do the choices of all judges on an instance vary?
Figure 5: Disagreement Analysis (Figures (a)(b)) and Mutual Agreement Heatmap (Figures (c)(d)(e)) of LLM judges on MTBench and DevBench. The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset. The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
H.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike" or consistent they are across a set of evaluations. We denote two judges agree on an instance if their choices are identical. Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances. Fig. 5(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain response, compared to the case when they prefer different responses. The "without tie" agreement heatmap of the twelve judges on MTBench is explored in Fig. 5(d).
The heatmaps reveal clear “familial patterns” in the choices of these LLM judges. For instance, the GPT-4, GPT-4-Turbo, and GPT-4o series exhibit high agreement on MTBench, achieving over 70% with ties included and over 85% without. GPT-3.5 models didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
For Claude-3 models, similar familial patterns could be observed. Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Opus, likely due to their similar model structure within the same series. Interestingly, Claude-3.5-Sonnet and Claude-3-Sonnet do not exhibit a significantly high agreement, indicating that the upgrade from series 3 to 3.5 considerably impacts their judging capabilities.
Gemini models exhibit rather low mutual agreement where "familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges. In particular, the LLM judges could be primarily grouped by their capabilities; when judging capabilities are comparable, models within the same family series share a higher mutual agreement than across families.
Identifying such groupings provides valuable insights, as comparisons between judges from different groups, both adept at assessing LLM-generated content, can reveal distinct position biases and enrich our understanding of this phenomenon.
H.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better response will have a disagreement of 0; an instance where judgments are widely varied will result in a high disagreement. For our study where twelve judges are investigated, the maximum disagreement of an MTBench is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig.5(a) and (b), respectively. From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks. These are likely easy-to-evaluate instances, and the reliability of LLM judgments is enhanced by majority voting. In comparison, the instances with the highest disagreement are likely the ones that are difficult to evaluate and where the position bias is most likely to occur. However, luckily, these instances are rare, occupying only 0.03% and 0.07% of MTBench and DevBench respectively. In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig.5(a)(b) exhibit a balanced distribution of instances with varied difficulty. This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
To summarize, the practical implications of the disagreement analysis are three-fold. First, it helps identify the instances that are difficult or trivial to judge, benefiting benchmark designs to control the difficulty of evaluation by managing the number of these instances across the dataset. Second, it assists in filtering out instances where majority voting of LLM evaluators are likely to offer reliable judgments without direct comparison with human-annotated evaluations, enhancing the scalability of LLM judges especially when human evaluations are costly. In other words, if one-shot judgments from only one LLM judge are not enough reliable, multiple capable LLMs and the majoring voting strategy could be employed to make the evaluation more convincing. Last but not least, disagreement analysis provides a convenient way to make the difficulty variance of instances varied across the dataset tangible. Since the difficulty of an evaluation instance is closely related to the quality gap between the two answers and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
Appendix I More Discussions on Answer Quality GAP
Intuitively, answer pairs with larger quality gaps are easier for LLM judges to evaluate, as one answer is clearly superior. Therefore, we expect LLM judges to exhibit weaker position bias in such cases. This motivates us to quantify the answer quality gap and investigate its impact on position bias. The quantification utilizes the overall win rate ($owr$) of a candidate model (i.e., answer provider) on the other to compute the quality gap. For a set of questions on a certain task and the corresponding responses answered by two models, an LLM judge’s evaluations would pick out the winner model. For the position consistent judgment instances, we can count the number of a model’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively. On the other hand, following the methodology introduced by [3], we consider the inconsistent judgment pairs as a tie for both models, whose counting number is denoted as $C_{I}$. This follows the intuition that it is difficult to strictly pick out a better one between two responses with similar quality, hence increasing the likelihood of position bias.
Therefore, considering both cases, we define the overall win rate ($owr$) of a model’s response over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs involving the two models in both original and swapped settings. Then the quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, as 0.5 is the exact tie case for the win rate ranging from 0 to 1. We prefer the overall win rate over the consistent win rate (calculated by $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ to mainly address the data sparsity issue. If the proportion of position inconsistent judgments becomes high, the available data becomes sparse and $n_{C}$ becomes small. In the extreme case, where all judgments are inconsistent, $n_{C}$ is 0 and the consistent win rate is not even computable, making the measurement inaccurate and impractical. In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie" and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
Appendix J More Results of Position Bias and Answer Quality Gap Measurement
As shown in Fig.4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness. In this session, we explore whether the discovery is consistent for each individual judge. Same as Section B.2, we apply the overall win rate to reflect the answer quality gap for visualization.
Figure 6: Positional consistency and positional fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
As shown in Fig.6 (a), the “parabolic shape" is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency" applies to all models. However, Fig.6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges. For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape", GPT-4 and GPT-3.5 present nearly linear curves. In other words, while the former models align with the general tendency that a larger answer quality gap improves positional fairness, the latter ones preserve fairness regardless of the answer quality gap. This further demonstrates the necessity to investigate positional fairness in addition to consistency when evaluating a judge model’s position bias.
Appendix K DevBench
This session includes a similar baseline comparison analysis on DevBench as on MTBench. As shown in Fig.7, position bias is judge-dependent and task-dependent on DevBench as well, as $PC$ and $PF$ vary significantly across judges and tasks. Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks. For instance, for architecture design evaluations, gpt-4-1106-preview, GPT-4o, and Gemini-1.5-pro all surpass GPT-4. Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations. However, GPT-4 is still the best-performing model on UML sequence evaluations, with only GPT-3.5-Turbo can achieve comparable performance regarding certain detailed metrics (e.g., interaction complexity). These discoveries, aligning with the findings on MTBench, further necessitate the need to consider the trade-offs between positional consistency and fairness when selecting the optimal judge model for certain tasks.
Figure 7: Baseline comparisons of judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be GPT-4 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
Appendix L Experiment Settings
This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study. We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
L.1 Judges, Models, and Tasks
Judge
In this study, we choose seven GPT, four Claude, and three Gemini models as the judges. The specific versions for API call are specified as follows: o1-mini-2024-09-12 for o1-mini, gpt-4o-2024-05-13 for GPT-4o, gpt-4-1106-preview for GPT-4-turbo, gpt-4-0613 for GPT-4, and gpt-3.5-turbo-1106 for GPT-3.5-turbo; claude-3-5-sonnet-20240620, claude-3-opus-20240229, claude-3-sonnet-20240229, and claude-3-haiku-20240307 for Claude series. The other model names and versions are as they are.
Model
The reference (or baseline) answer-generating models are vicuna-13b-v1.3 for MTBench and human for DevBench. They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations. The other models that are compared to the reference models, namely “Model" in our context, are listed as follows.
  • MTBench (30): alpaca-13b, baize-v2-13b, chatglm-6b, claude-instant-v1, claude-v1, dolly-v2-12b, falcon-40b-instruct, fastchat-t5-3b, gpt-3.5-turbo, gpt-4, gpt4all-13b-snoozy, guanaco-33b, guanaco-65b, h2ogpt-oasst-open-llama-13b, koala-13b, llama-13b, mpt-30b-chat, mpt-30b-instruct, mpt-7b-chat, nous-hermes-13b, oasst-sft-4-pythia-12b, oasst-sft-7-llama-30b, palm-2-chat-bison-001, rwkv-4-raven-14b, stablelm-tuned-alpha-7b, tulu-30b, vicuna-33b-v1.3, vicuna-7b-v1.3, wizardlm-13b, wizardlm-30b
  • DevBench (10): codellama-7b-instruct, codellama-13b-instruct, codellama-34b-instruct, deepseek-coder-1.3b-instruct, deepseek-coder-6.7b-instruct, deepseek-coder-33b-instruct, gpt-3.5-turbo-1106, gpt-4-0125-preview, gpt-4-0613, gpt-4-1106-preview
The model names are exactly what MTBench [3] and DevBench [4] used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
Task
For tasks, we also follow the original studies of these two benchmarks, except for DevBench we separate the gerenal metrics into detailed ones and considered them as different tasks. In this sense, our study experiments on the following tasks to provide a comprehensive study on the positon bias of LLM-as-a-Judge:
  • MTBench (8): coding, extraction, humanities, math, reasoning, roleplay, stem, and writing.
  • Devbench (14):
    • UML class (4): cohesion_and_decoupling, complexity, practicability, and faithfulness
    • UML sequence (5): cohesion_and_decoupling, interaction_complexity, practicability, uniformity_and_integration, and faithfulness
    • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
L.2 Prompt Settings
We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.
Though written differently, these prompts all share same key components:
  • A system prompt explaining the judging task and the role the LLM should be playing.
  • Emphasized “should" and“shouldn’t"s.
  • A prompt structure with placeholders for specific questions and model answers
  • A specified output format for later judgment extraction
  • Chain-of-Thought [53] prompts requiring the LLM judge to provide reasons for its judgment
The detailed prompt templates are specified below.