### v7

# Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge

Abstract
LLM-as-a-Judge has emerged as a promising alternative to human evaluators across various tasks, yet inherent biases—particularly position bias, the tendency to favor solutions based on their position within the prompt—compromise its reliability. This study investigates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness. Our experiments, involving 12 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 100,000 evaluation instances. We identify Judge-Level, Candidate-Level, and Task-Level factors contributing to bias. The findings confirm that position bias is not due to random chance and varies significantly across judges and tasks. While position bias is weakly influenced by the length of prompt components, it is strongly affected by the quality gap between solutions. Our agreement and disagreement analysis among judges further provides insights into the distribution of judging difficulty across the dataset, and highlights the potential for dataset modifications.
Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge
Lin Shi, Chiyu Ma, Wenhua Liang, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
1 Introduction
Figure 1: Overview of our experiment settings: (a) Position bias is observed when LLM judges consistently favor a specific position rather than evaluating the content, with repeated trials ensuring the deviations are not due to random variations. (b) Preference fairness is defined and measured through the distribution of choice pairs to assess the fairness of judgments. (c) The settings are extended from pairwise comparisons to list-wise comparisons, involving evaluations of more than two candidate models.
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations Zeng et al. (2023); Karpinska et al. (2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments Zheng et al. (2024b); Li et al. (2024a); Zhu et al. (2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1 (a). This bias has been prevalent in all types of LLM judges Qin et al. (2024); Li et al. (2023d), raising concerns about their reliability. Previous studies Zheng et al. (2024a, b); Zeng et al. (2023) have identified position bias alongside other biases and assessed its impact. Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al. (2024), increased complexity Li et al. (2024b); Khan et al. (2024); Chua et al. (2024), the introduction of new biases Ohi et al. (2024), or inconsistent effectiveness Gallegos et al. (2024). These persistent challenges highlight the limitations of current approaches and point to a possible incomplete understanding of position bias. Although numerous studies have evaluated position bias in LLM-as-a-Judge settings, the underlying factors contributing to this phenomenon remain insufficiently explored.
In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge. While evaluating Position Consistency, one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability. Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences. Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
To investigate the underlying factors contributing to position bias, we categorized these factors into three levels: Judge-Level, Candidate-Level, and Task-Level. Our experiments are primarily conducted on pairwise comparisons, as LLM judges demonstrate superior performance in this setting. We further extend our study to more complicated list-wise comparison settings, involving evaluations of more than two candidate models by LLM judges. Our findings reveal several key insights: 1. The position bias of capable LLM judges is not a result of random variations. 2. There is a high volatility in the direction of preference, even within the same LLM judge when applied to different tasks. 3. Differences in answer quality among candidate models significantly influence position consistency. 4. Position bias is very weakly correlated with the length of prompts generated by the candidate models.
Building on these findings, we conduct an agreement analysis among the LLM judges. The results reveal that, although measures of position consistency may appear similar in general, judgments on specific instances vary significantly among LLM judges, even when they demonstrate comparable capabilities. Instances where numerous LLMs agree are generally easier to judge, whereas instances with disagreements are more challenging to evaluate and more prone to position bias. This analysis provides insights into the distribution of judging difficulty across the dataset and highlights the potential for dataset modifications by incorporating more instances that are either easier or more difficult to judge. Future work could explore how to measure the likelihood of position bias arise from the datasets by identifying and quantifying such hard-to-judge instances before implementing LLM judges.
2 Evaluation Settings & Definitions
We begin by outlining the settings for pairwise and list-wise comparisons employed in our experiments for LLM-as-a-Judge. Following this, we define the three metrics used in our evaluation: Position Consistency (PC), Preference Fairness (PF), and Repetition Stability (RS). Finally, we provide a detailed description of the factors we found that are related to position bias at the Judge-Level, Candidate-Level, and Task-Level.
2.1 Pairwise & List-wise Comparison
Pairwise Comparison: In the context of pairwise comparison, LLM judges are tasked with selecting the better solution provided by two candidate models in response to a given task question. As shown in Fig. 1 (a), the system prompt, option choices, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the better solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair. If the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy (e.g. always choose $\{A\}$) or recency (e.g. always choose $\{B\}$). Example of measuring preference fairness with specific choice pairs is shown in Fig. 1 (b). To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts judges to choosing between two options, labeled A for the first candidate and B for the second. Three-Option mode introduces an additional choice, C, allowing judges to indicate a tie if neither solution is preferable, as illustrated in Fig. 1 (a). These option modes were explicitly specified in the system prompts to ensure clear guidance for the decision-making process of the LLM judges.
List-wise Comparison: Unlike pairwise settings, where LLM judges select the superior solution from two candidates, list-wise comparative approaches involve evaluating three or more candidates simultaneously, as shown in Fig. 1 (c). For efficiency, we prompt judges to select the best candidate rather than ranking the entire list. The “swapped setting" used in pairwise evaluations is generalized to order permutations for list-wise judgments, ensuring that each candidate appears in every possible position exactly once. For a list of $p$ candidates, this results in $p$ permutations. In the $i^{th}$ permutation, the $i^{th}$ candidate is set to appear in the first position. Additionally, an option is provided to account for ties, allowing judges to indicate if there is no certainly superior solution.
2.2 Evaluation Metrics
In our study, we first verify if the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows:
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or by random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC$ $=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{k\in S}\left\{|C_{k}^{j}|\right\},$ (1)
where $S=\{A,B,C,\dots\}$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries. The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is permuted. It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the judge model prefers the same winning solution across permutations. Formally, it is calculated as
$PC$ $=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{1}^{j},\dots,C_{p}^{j},)\in V\}},$ (2)
where $V$ is the set of choices that correspond to position consistency, and $(C_{1}^{j},\dots,C_{p}^{j})$ denotes the judgment series for the $j^{th}$ query when there are $p$ candidate solutions in the list, and $n$ represents the number of prompt series. An example of such series of choices under pairwise comparison setting can be found in Fig. 1 (b). This formula aims to provide a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
Preference Fairness ($PF$) measures the extent to which judge models favor certain solution positions. In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second candidate model" used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 1 (b). Previous studies proposed two common ways to measure the preference fairness. One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach overlooks the fact that “position consistent judgments are also preference fair”, which leads to overly penalizing highly consistent LLM-judges.
To overcome these limitations, we introduce a more granular and scalable measurement that combines the strengths of both methods, to assess preference fairness. The $PF$ score is formally calculated by
$PF$ $=\frac{PF_{\text{raw}}-S^{-}_{\min}}{S^{+}_{\max}-S^{-}_{\min}}\times 2-1,$ (3)
$PF_{\text{raw}}$ $=(rcn\times irr)-(pcn\times ipr).$
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scale ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. The $PF$ score is interpreted as follows:
$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
To extend this metric to list-wise comparisons, we employed a ‘one vs. all’ approach, defining primacy preference as favoring the first candidate solution while classifying all others as recency-preferred. This straightforward extension of the $PF$ computation maintains consistency with pairwise setups. By providing a single and comprehensive metric that applies to all evaluation instances and list-wise settings, our proposed $PF$ score ensures sensitivity across datasets, regardless of variations in the number of questions or instances, offering a significant improvement over previous methods.
2.3 Factors Affecting Position Bias
Factor Judge-level Candidate-level Task-level
Familial Property $\checkmark$* $\times$ $\times$
Answer Quality Gap $\times$ $\checkmark$* $\times$
Task Input Length $\times$ $\times$ $\checkmark$
Task Output Length $\times$ $\checkmark$* $\checkmark$*
Prompt Length $\times$ $\checkmark$ $\checkmark$
Table 1: Factors influencing position bias. Significant factors, identified via bidirectional stepwise regression, are marked with * and highlighted in red based on empirical findings. Task Input refers to the question itself, while Task Output denotes the candidate model’s answers, serving as both Candidate-level and Task-level factors. Prompt includes the full query presented to judges: Task Input, Task Output, and system prompts.
To investigate the factors influencing position bias in LLM judges, we categorized these factors into three groups: Judge-level, Candidate-level, and Task-level factors. Each group includes specific factors, that we hypothesize, may impact position bias, which we explore through a series of experiments. Table 1 lists the five factors we analyzed in this study.
Among the influencing factors, we selected “familial property” for Judge-level factors, as it reflects similar model sizes and training specifics, which are often proprietary and not publicly accessible. The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version. More details and discussions about the familial property can be found in Appendix Sec. C.
Answer quality gap, based on our findings, is highly influential yet under-explored in prior research. We define the quality of a candidate’s solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
Following this assumption, we measure the answer quality gap by the win rates of candidates over an expected baseline on a set of tasks and questions. However, if position bias occurs, the winner may be inconsistent when the order of candidate solutions is permuted in the query. Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen across all permutations (termed “consistent wins”), cases where there is no certain winner (termed “consistent ties”), and cases where different winners are selected after the solutions are permuted in the queries (termed “inconsistent judgment series"). We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment series ($C_{I}$), respectively. Inspired by Zheng et al., we count inconsistent judgment pairs as ties for all candidate models, which is later calculated as a down-scaled win rate depending on the number of candidate models.
To calculate the win rates of candidate models for all three cases, we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{p}(C_{t}+C_{I})}]$, where we have $p$ candidates in the list and $n$ judgment series. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-1/p|$, where $1/p$ is the expected baseline when all judgments are “ties”. In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment series) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
3 Experiment
3.1 Experiment Settings
In this study, we evaluated twelve up-to-date close-sourced models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework. Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks. However, due to their limited context window size and failure to follow the output format, which led to suboptimal performance on our tasks (as shown in Table 2), we focused our investigation exclusively on more powerful closed-source models. We included the results of open-source models for completeness. More discussions can be found in Appendix.Sec. D
We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks. For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses. MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges. We adopted Two-option mode for MTBench, and Three-option mode for DevBench. For list-wise experiments, we randomly sampled 9 models to form three triple-candidate lists and evaluated four representative judges on MTBench. The prompt templates we used are identical to those in the benchmarks for pairwise comparisons, with minor modifications to accommodate list-wise evaluations. More details about the models, tasks, and prompts can be found in Appendix. Sec. G.
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials. This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench. The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, more than 100,000 evaluation instances were analyzed in this study.
To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks. We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively. Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score. This process involves both forward selection and backward elimination, with each "step" testing whether including or excluding a variable improves the model’s AIC value. Further details about the process can be found in Appendix. Sec. F.
3.2 Empirical Results
The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not due to random variations. Fig. 2 (a)(b) demonstrate that position bias varies by judges and tasks significantly. Fig. 2 (c) explores the correlation between the metrics $PC$ and $PF$. Fig. (d) to (f) further investigate the impact of the answer quality gap on position bias. These analyses were conducted by considering all judges together on MTBench. The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E.
Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables. Similarly, these factors also contribute significantly to Preference Fairness. Additionally, we found that average output length is a statistically significant predictor of $PF$. This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred. Quantitative results and more discussions can be found in Appendix. Sec. F.
Judge MTBench Pairwise DevBench Pairwise MTBench List-wise
$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error $PC$ $PF$ Error
Claude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 0.67 $\pm$ 0.19 0.17 $\pm$ 0.19 0.00
Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00
Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00
Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00
Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 0.55 $\pm$ 0.20 0.33 $\pm$ 0.18 0.00
Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 0.68 $\pm$ 0.22 0.18 $\pm$ 0.22 0.00
o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 0.34 $\pm$ 0.17 -0.05 $\pm$ 0.30 0.12
Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold. Errors arise from judgment failures (e.g., exceeding context window, not following output format). High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data. Open-source models (e.g., Llama) are excluded due to high error rates with our lengthy instances (See Appendix.Sec.D for more discussions). List-wise evaluation is conducted on four representative judges from each family to validate scalability.
Figure 2: Judge performances on MTBench. Fig. (a)(b) are the radar charts for the $PC$ comparison by family, judge, and task Fig. (c) to (f) investigate the impact of task output length and answer quality gap on position bias. Fig. (a)(c)(d)(e) are for pairwise comparative settings, while Fig. (b)(f) are obtained under list-wise evaluations.
4 Main Findings
Position Bias of Capable Judges are not Mere Random Variations: As shown in Table 2, the capable judges on the benchmark tasks, supported by minimal "Error" rates, generally exhibit $RS$ values above 0.85. The most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks. These results confirm that LLM judgments, and the resulting position bias, are not due to random variations. This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
Position Bias Varies by Judge & Task: As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig. 2 (a) for $PC$ comparisons. Moreover, certain models achieve comparable or superior performances than GPT-4 on certain tasks. For example, according to Fig. 2 (a), for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges. Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it could be employed as a cost-effective alternative to coding evaluations.
Additionally, significant variations of $PF$ across tasks are observed. The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others. Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks. Even for judges that are recency-preferred across all tasks (e.g., Claude-3’s and Gemini-pro’s), the extent of biased preference, as reflected by $PF$ values, varies by task.
Moreover, a high position consistency does not guarantee fairness. For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively. In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
In list-wise comparisons, similar variations by judge and task in position bias were observed, as shown in Fig. 2 (b). Furthermore, Table 2 highlights that more capable models, such as GPT-4o and Claude-3.5-Sonnet, maintain high consistency when transitioning from pairwise to list-wise evaluations, while less capable models, such as GPT-3.5-Turbo, exhibit greater sensitivity to the increased number of candidates in list-wise tasks.
Therefore, the position bias of LLM judges is judge-dependent and task-dependent. This observation is further confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$. In practice, to achieve the most reliable evaluations, one needs to select task-specific LLMs whose consistency and fairness are in balance. Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec. E.2.
Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate ($owr$). Therefore, $\delta_{q}$ increases as $owr$ extends from baseline to 0 or 1. Fig. 2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answers with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. E.
Similarly, as shown in Fig. 2 (d), judgments generally become more preference fair as $\delta_{q}$ increases. However, the extent is not as significant as for $PC$. Also, the relationship varies by judge, as some LLMs maintain preference fairness regardless of $\delta_{q}$. For example, as shown in Appendix.Fig. 6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench. These observations align with the right-arrow shape as demonstrated in Fig. 2 (c), where there is a general trend that judgments become preference fairer as position consistency increases. It also justifies the reasonableness of our quantification of preference fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
Together, we conclude that as the answer quality gap enlarges, judges generally become more position consistent and preference fair according to the regression curves. However, exceptions are common, as shown by the individual scatter points of these figures. This indicates that though the answer quality gap significantly influences the position bias of LLMs, other factors also play important roles. Therefore, built on our findings, future studies may have better control over the answer quality gap when evaluating LLM judges, exploring other impacting factors on position bias, and seeking potential mitigation strategies.
Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length). By stepwise regression, we discovered that average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table. 4. In other words, there is a very weak relationship between the lengths of prompt components and position bias.
LLM Agreement Analysis: We complement our investigation of position bias with an agreement and disagreement analysis among LLM judges. Rather than focusing exclusively on overall consistency or fairness, we examine how judges converge and diverge in their assessments of individual instances. Agreement analysis quantifies the percentage of instances where two judges mutually agree on the outcome. Disagreement analysis counts the number of choices deviating from the mode for each instance among all judges. This further complies a “distribution of disagreement” across the dataset.
Our findings reveal that, despite exhibiting similar overall $PC$ and $PF$ scores, judges vary significantly in their judgments on individual instance. Disagreement analysis, in particular, highlights instances where consensus is either easily or difficultly achieved, reflecting the inherent complexity of the judgment task. For example, as shown in Fig. 3, more than half of the dataset can be considered relatively easy to judge, as over 75% of all 12 judges agree with each other on these instances. Conversely, fewer than 4% of instances represent the likely hard-to-judge cases where a majority of judges fail to reach consensus.
Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among judges does not exceed a specific threshold.
Based on our observations of answer quality gaps and LLM agreement/disagreement patterns, this study offers practical insights for designing evaluator benchmarks that account for the varying difficulty levels of judgment tasks. Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of judges. Further discussions and analyses can be found in Appendix Sec. C.
5 Conclusion
In conclusion, this paper provides an in-depth analysis of position bias in LLM judges, a critical challenge in automated evaluation. Using metrics such as repetition stability, position consistency, and preference fairness, we identify significant variations in position bias across judges and tasks, consistent across pairwise and list-wise comparison settings. Our findings show that position bias is weakly influenced by prompt length but strongly impacted by the quality gap between solutions. Furthermore, agreement and disagreement analysis highlights variability in judgment reliability, providing valuable insights for refining benchmarks. This study enhances understanding of position bias and contributes to the development of fairer and more reliable LLM evaluation frameworks.
6 Limitations
Despite proposing scalable metrics and investigating key factors influencing position bias, our study has several limitations. First, due to computational constraints, we evaluated only twelve commercial LLM judges for pairwise settings and four for list-wise settings, limited to 3-candidate lists, using two benchmarks. Open-source models, such as Llama 7B-13B, struggled to provide reliable judgments, restricting their inclusion. Additionally, while we tested original benchmark prompt templates, exploring variations in prompting techniques could provide further insights. Future work could expand this study to include more models, tasks, prompting strategies, and larger list-wise candidate pools to validate and generalize our findings.
Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size. Instead, we approximated these factors by grouping models by family properties. Open-source models with accessible architectural details could allow deeper analysis of these influences. Additionally, our analyses rely on completed judgments, such as measuring the answer quality gap and task difficulty after evaluations. Future work could explore methods to estimate or control these factors pre-judgment, reducing computational costs and enabling proactive mitigation strategies.
Lastly, our focus was on understanding position bias rather than mitigating it. While our findings provide a foundation for effective mitigation, further research is needed to address issues like maintaining consistency and fairness when answer quality gaps are minimal, where position bias is most pronounced. Multivariate analyses exploring interactions between factors like prompt length, task complexity, and answer quality gaps could also yield deeper insights and enhance mitigation approaches.
Appendix Table of Contents
[…]
[appendices]0
Appendix A Reproducibility
Our experiments were conducted primarily using API access, with a total cost of approximately 3,000 USD.
The code repository for reproducing our results is available at: https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
Appendix B Related Work
B.1 LLM-as-a-Judge
Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments. The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (Chiang and Lee, 2023a), adversarial attacks (Chiang and Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi and Federmann, 2023), and instruction following (Zeng et al., 2023), where models are tasked with scoring or ranking outputs. Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness, even in the most effective pairwise comparative settings.
B.2 Position Bias
However, position bias is a complex problem that is challenging to deal with or solve. The naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited. Therefore, researchers proposed ‘inconsistency-as-a-tie’ for both models in pairwise comparative settings to consider all judgments for further analysis (Zheng et al., 2024b; Li et al., 2023c). This approach, while practically useful for evaluations, does not mitigate position bias.
Given the significance of position bias, researchers have developed more sophisticated approaches to mitigate it, including bootstrapping (Hou et al., 2024), split-and-merge techniques (Li et al., 2023d), and multi-agent discussions (Li et al., 2023c; Khan et al., 2024). While these methods demonstrate potential, they are often costly, time-consuming, or insufficient. For example, multi-agent discussions and iterative reviews require extensive computational resources and human-like deliberation, while simpler techniques, such as order permutation (Wang et al., 2023; Zheng et al., 2024a) and batch calibration (Zhou et al., 2024), have limited effectiveness in addressing bias holistically. Furthermore, calibration methods like ensembling (Li et al., 2023a) or context decomposition (Li et al., 2023d) are often constrained by the inadequacy of supervised data to align LLM evaluators effectively (Liu et al., 2024). This highlights that addressing this crucial bias is complex and challenging.
The limitations of existing approaches underscore a critical gap in our understanding of position bias itself. A lack of clarity on the factors influencing this bias and their quantitative impacts hampers the development of effective solutions. For instance, while the PORTIA approach (Li et al., 2023d) demonstrates significant improvements in consistency for GPT-4 (from 93.44% to 97.03%), this success may be overestimated due to the selection of tasks with large answer quality gaps. As our study reveals, such gaps inherently reduce the likelihood of position bias, making these cases easier to calibrate. This example underscores the necessity of first comprehensively understanding the underlying factors contributing to position bias before proposing or refining mitigation strategies. Without this foundational understanding, the efficacy of current and future methods remains uncertain and potentially misjudged.
B.3 Research Gaps
To enhance the understanding of position bias, our study builds on and extends prior work through the following dimensions.
Repetition Stability: Previous studies on LLM-as-a-Judge often implicitly assume that single-instance judgments are sufficient to reflect the evaluating capabilities of judge models, thereby attributing observed biases entirely to systematic tendencies rather than random variations. These studies typically proceed directly to analyses on position bias or others without such verification. However, if judgments vary by chance or lack consistency across repetitions, the reliability and accuracy of downstream analyses are compromised. To avoid this issue, Chen et al. (2024) repeated evaluations multiple times and adopted the mode judgment as the representative, an approach that, while effective, is computationally expensive. Addressing this gap, our work introduces a method or metric for systematically evaluating the stability of LLM judgments across repetitions. This ensures that observed biases are systematic rather than artifacts of variability. By performing this validation on a sample dataset for cost-efficiency, researchers can establish a robust foundation for subsequent evaluations and analyses, ensuring the reliability of their conclusions.
Preference Fairness: Prior studies on position bias often focus on the proportion of position consistent judgments, referred to as “position consistency”. Only a few have explored the preference directions of LLM evaluators in addition to consistency. However, traditional measurements either overly penalize highly consistent judges (Li et al., 2023c; Liusie et al., 2024) or rely on metrics whose sensitivity is highly dependent on dataset size (Zheng et al., 2024b; Zhu et al., 2023). To address these limitations, we proposed a refined metric, preference fairness ($PF$), which integrates the strengths of both approaches. This single, comprehensive metric ensures sensitivity across datasets, remains independent of the number of questions and tasks, and represents a significant improvement over previous metrics.
Answer Quality Gap: Intuitively, candidate solutions with significant quality disparities are easier to judge, likely resulting in reduced position bias. This suggests that the quality gap among answers is a potentially significant factor influencing position bias. However, this factor has been largely overlooked in prior works, both in evaluations and in the development of mitigation strategies, underscoring our limited understanding of position bias. While Wang et al. (2023) explored the answer quality gap in score-based LLM-as-a-Judge evaluations using judgment scores, quantifying this variable in comparative settings (pairwise or list-wise) remains an open research problem. To address this gap, we propose the use of win rates (Zheng et al., 2024b) of candidate solutions against an expected baseline to measure the answer quality gap. This approach offers a practical quantitative metric for subsequent analysis and deepens our understanding of the factors driving position bias.
List-wise Evaluations: Although LLM-as-a-Judge has been categorized into pointwise, pairwise, and list-wise paradigms (Qin et al., 2024), most research has focused on pairwise comparative assessments due to their demonstrated effectiveness (Zheng et al., 2024b; Liusie et al., 2024). However, this leaves a significant gap in measuring and understanding the position bias in list-wise settings, where LLMs are prompted to evaluate three or more candidate solutions simultaneously. With advancements in model capabilities, list-wise evaluations present a potentially more cost-effective alternative to pairwise approaches, requiring fewer comparisons while maintaining consistency and fairness. Our work bridges this gap by extending position bias assessments from pairwise to list-wise settings, enabling a more comprehensive evaluation of LLM-as-a-Judge across varied configurations.
These identified gaps - spanning from the validation of judgments to the evaluation of list-wise settings - highlight the need for a more comprehensive framework to evaluate and understand position bias. By addressing these dimensions collectively, our work provides a foundation for a deeper understanding of the consistency, fairness, and reliability of LLM evaluators.
Appendix C LLM Agreement Analysis
Besides the exploration of position bias with a broad lens by average $PC$ and $PF$, instance-wise agreement between LLM judges is also insightful. Even two judges with the same $PC$ and $PF$ scores may not reach consensus on each instance. Therefore, this session investigates (1) what percentage of a set of evaluations do two LLM judges agree on each other? (2) how do the choices of all judges on an instance vary?
Figure 4: Disagreement Analysis and Mutual Agreement Heatmap of LLM judges on MTBench and DevBench. The blue bars of Figures (a) and (b) represent the number of instances with certain disagreement values, and the red cumulative curve displays the cumulative count and corresponding percentage over the benchmark dataset. The heat maps of Figures (c)(d)(e) marked higher mutual agreement values with a brighter color.
C.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations. We denote two judges agree on an instance if their judgment choices are identical. Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances. Fig. 4(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions. The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig. 4(d).
The heatmaps reveal clear “familial patterns” in the judgment choices of these LLM judges. For instance, the GPT-4, GPT-4-Turbo, and GPT-4o series exhibit high agreement on MTBench, achieving over 70% with ties included and over 85% without. GPT-3.5-Turbo didn’t agree with the GPT-4 series and o1-mini for around 40% of the instances, indicating that they are considerably different in judging capabilities.
For Claude-3 models, similar familial patterns could be observed. Claude-3-Opus highly agrees with Claude-3.5-Sonnet, probably due to their similar capabilities, while it also highly agrees with Claude-3-Sonnet, likely due to their similar model structure within the same series. Interestingly, Claude-3.5-Sonnet and Claude-3-Sonnet do not exhibit a significantly high agreement, indicating that the upgrade from series 3 to 3.5 considerably impacts their judging capabilities.
Gemini models exhibit rather low mutual agreement and “familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
These patterns suggest that familial similarities, possibly stemming from analogous model sizes, training data, and strategies, influence the positional preferences of these judges. In particular, the LLM judges could be primarily grouped by their capabilities; when judging capabilities are comparable, models within the same family series share a higher mutual agreement than across families.
Identifying such groupings provides valuable insights, as comparisons between judges from different groups, both adept at assessing LLM-generated content, can reveal distinct position biases and enrich our understanding of this phenomenon.
C.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement. For our study where twelve judges are investigated, the maximum disagreement of an MTBench instance is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
The distributions of instances with different disagreement values on MTBench and DevBench are shown in Fig. 4(a) and Fig. 4(b), respectively. From our disagreement analysis, at least 75% of the judges reached a choice consensus on more than half of the instances on both benchmarks. These are likely easy-to-evaluate instances, and the reliability of LLM judgments is enhanced by majority voting. In comparison, the instances with the highest disagreement are likely the ones that are difficult to evaluate and where the position bias is most likely to occur. However, luckily, these instances are rare, occupying only 0.03% and 0.07% of MTBench and DevBench respectively. In other words, majority voting of multiple capable LLM judges could be practically useful for over 99% of evaluation instances on both benchmarks.
Moreover, if we roughly consider the disagreement value of instances as their difficulty for judging, then Fig. 4(a) and Fig. 4(b) exhibit a balanced distribution of instances with varied difficulty. This is because, except for the instances with the highest disagreement, the numbers of other instances with varied disagreement do not vary significantly, indicating a smoothly increasing difficulty curve across the benchmark datasets.
To summarize, the practical implications of the disagreement analysis are three-fold. First, it helps identify the instances that are difficult or trivial to judge, benefiting benchmark designs to control the difficulty of evaluation by managing the number of these instances across the dataset. Second, it assists in filtering out instances where majority voting of LLM evaluators are likely to offer reliable judgments without direct comparison with human-annotated evaluations, enhancing the scalability of LLM judges especially when human evaluations are costly. In other words, if one-shot judgments from only one LLM judge are not enough reliable, multiple capable LLMs and the majoring voting strategy could be employed to make the evaluation more convincing. Last but not least, disagreement analysis provides a convenient way to make the difficulty variance of instances varied across the dataset tangible. Since the difficulty of an evaluation instance is closely related to the quality gap between the two solutions and hence position bias, the investigation of the instances where most judges particularly disagree with one another could provide more insights and inspiration for future benchmark designs and potential mitigation strategies for position bias.
Appendix D Open-source Model Discussion
Due to computational constraints, we conducted limited experiments on open-source models, specifically Llama 7B-13B, as representatives of smaller-scale models. However, as shown in Table 2, these models struggled to perform effectively on our evaluation tasks. A primary issue was their limited context window, which proved insufficient for benchmarks like DevBench, where questions often include extensive instructions, such as full product requirement documents. This context limitation caused frequent truncation of either the prompt or the candidate responses, leading to incomplete or erroneous outputs and significantly higher error rates.
Figure 5: Llama-13B output example: Despite providing analysis and a choice, the output format does not match the specified “[[choice]]" pattern, causing extraction failures.
Another key factor affecting performance was mismatched output formatting. MTBench requires responses in the format [[choice]] (e.g., [[A]]), but Llama-13B, as shown in Fig. 5, frequently produced outputs such as [A], which could not be parsed correctly by the regular expression used for choice extraction. Interestingly, Llama-7B outperformed Llama-13B on MTBench, likely due to its simpler, shorter responses that inadvertently aligned better with the expected format. These errors illustrate the challenges these models faced in meeting benchmark-specific requirements, highlighting nuances in their evaluation results. Considering practicality and effectiveness, we only focus on more capable commercial judge models.
Appendix E More Results of Position Bias and Answer Quality Gap Measurement
E.1 MTBench
As shown in Fig. 2(c), considering all judges together, a larger answer quality gap generally leads to better position consistency and preference fairness. In this session, we explore whether the discovery is consistent for each individual judge. Same as Section 2.3, we apply the overall win rate to reflect the answer quality gap for visualization.
Figure 6: Position Consistency and Preference Fairness vs. overall win rate for each judge on MTBench. Figure (a) refers to the relationship investigation of $PC$ and figure (b) for $PF$.
As shown in Fig. 6 (a), the “parabolic shape" is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher position consistency applies to all models. However, Fig. 6 (b) reveals that preference fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges. For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape", GPT-4 and GPT-3.5 present nearly linear curves. In other words, while the former models align with the general tendency that a larger answer quality gap improves preference fairness, the latter ones preserve fairness regardless of the answer quality gap. This further demonstrates the necessity to investigate preference fairness in addition to consistency when evaluating a judge model’s position bias.
E.2 DevBench
This session includes a similar baseline comparison analysis on DevBench as on MTBench. As shown in Fig.7, position bias is judge-dependent and task-dependent on DevBench as well, as $PC$ and $PF$ vary significantly across judges and tasks. Similarly, although GPT-4 stands as the baseline model with a generally high $PC$ across tasks, certain models achieve comparable or superior performances on certain tasks. For instance, for architecture design evaluations, GPT-4-Turbo, GPT-4o, and Gemini-1.5-pro all surpass GPT-4. Gemini-1.5-pro is especially outstanding, also exceeding GPT-4 in uml class evaluations. However, GPT-4 is still the best-performing model on UML sequence evaluations, with only GPT-3.5-Turbo can achieve comparable performance regarding certain detailed metrics (e.g., interaction complexity). These discoveries, aligning with the findings on MTBench, further necessitate the need to consider the trade-offs between positional consistency and fairness when selecting the optimal judge model for certain tasks.
Figure 7: Baseline comparisons of judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be GPT-4 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
Appendix F Variable Selection and Tests
F.1 Bidirectional Stepwise Regression with AIC
Bidirectional stepwise regression is a combination of forward selection and backward elimination techniques. It iteratively refines the model by adding or removing predictors based on a statistical criterion—commonly the Akaike Information Criterion (AIC). The objective is to select a model that balances goodness of fit and complexity, aiming for the lowest AIC value.
The AIC is given by:
$\text{AIC}=2k-2\log(L),$ (4)
where $L$ is the likelihood of the model and $k$ is the number of parameters in the model, including the error variance $\sigma^{2}$. For a linear regression model with independent and identically distributed (iid) errors, $N(0,\sigma^{2})$, fitted to $n$ observations, the log-likelihood can be written as:
$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$
where $\hat{e}_{i}$ is the residual for the $i$th observation, and $\sigma^{2}$ is the variance of the errors. The AIC, in this context, becomes:
$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$
This form of the AIC balances the goodness of fit (as reflected by the residual sum of squares) and model complexity (as represented by $k$).
The operation of Bidirectional stepwise regression starts with either no predictors (forward selection) or all predictors (backward elimination), where the model iteratively adds or removes variables. Each step evaluates the impact on the AIC score. In forward selection, variables are added one by one, starting from the null model, such that the addition of each variable results in the largest decrease in AIC. In backward elimination, all variables are included in the model initially, and variables are removed one at a time, with the variable whose removal causes the smallest increase in AIC being dropped.
At each iteration, the change in AIC is computed as $\Delta\text{AIC}=\text{AIC}_{\text{new}}-\text{AIC}_{\text{current}}$, where $\text{AIC}_{\text{new}}$ refers to the AIC after adding or removing a variable, and $\text{AIC}_{\text{current}}$ is the AIC of the current model. If $\Delta\text{AIC}<0$, the model is improved by the addition or removal of the variable. The process terminates when neither adding nor removing variables results in a lower AIC, signifying that the most parsimonious model, based on AIC, has been reached.
F.2 Test results
We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias. Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively. Table 5, 6 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively. Table 7, 8 serves for DevBench, and Table 3, 4 is conducted on the integrated set of both benchmarks. The impact of variables on the model is ranked from highest to lowest, from bottom to top. Removed variables listed as None indicate the full model at this given step.
Through benchmark testing, we verified that LLM judges, task categories, and the answer quality gap significantly contribute to position bias in terms of both position consistency and preference fairness. These findings align with our empirical results, showing that position bias varies notably by judge and task, with the answer quality gap being a key influencing factor. The extent of this impact is reflected by the magnitude of change in AIC when the given variable is removed. It is worth noting that while task output length remains a significant predictor for $PF$ and $PC$ in both benchmarks, the change in AIC magnitude after removing this variable is very minimal. This is consistent across both benchmarks individually and combined. We therefore conclude that, although position bias is influenced by task output length, this dependency is minimal.
Removed Variables DF Sum of Sq RSS AIC
None 163.75 -18370
Task 20 2.832 166.59 -18319
Candidate 38 4.472 168.23 -18303
Quality gap 1 21.953 185.71 -17703
Judge 13 55.417 219.17 -16846
Table 3: Final results of stepwise model selection for both benchmarks: Position Consistency
Removed Variables DF Sum of Sq RSS AIC
None 254.28 -16103
Task output length 1 0.836 255.12 -16088
Quality gap 1 11.339 265.62 -15873
Task 21 16.177 270.46 -15817
Judge 13 82.069 336.35 -14641
Table 4: Final results of stepwise model selection for both benchmarks: Preference Fairness
Removed Variables DF Sum of Sq RSS AIC
None 61.974 -13312
Task output length 1 0.0553 62.029 -13311
Candidate 29 1.6474 63.621 -13282
Task 7 1.5304 63.504 -13244
Judge 13 15.3637 77.338 -12594
Quality gap 1 15.6206 77.594 -12559
Table 5: Final results of stepwise model selection for MTBench: Position Consistency
Removed Variables DF Sum of Sq RSS AIC
None 129.00 -10909.2
Quality gap 1 1.931 130.93 -10861.3
Task 7 9.295 138.29 -10689.4
Judge 13 58.847 187.85 -9672.5
Table 6: Final results of stepwise model selection for MTBench: Preference Fairness
Removed Variables DF Sum of Sq RSS AIC
None 55.382 -6940.2
Task output length 1 0.257 55.638 -6933.2
Candidate 9 1.514 56.896 -6905.4
Quality gap 1 13.128 68.510 -6525.3
Judge 13 84.760 140.141 -5146.6
Table 7: Final results of stepwise model selection for DevBench: Position Consistency
Removed Variables DF Sum of Sq RSS AIC
None 60.104 -6753.9
Task output length 1 0.061 60.165 -6753.9
Candidate 9 0.731 60.834 -6748.2
Task 13 1.305 61.408 -6737.8
Quality gap 1 1.783 61.886 -6698.6
Judge 13 80.875 140.979 -5108.9
Table 8: Final results of stepwise model selection for DevBench: Preference Fairness
Appendix G Experiment Settings
This session specifies more detailed information about the judges, answer-generating models, tasks, and prompt templates used in this study. We choose to evaluate MTBench and DevBench for the following reasons: (1) all necessary information about the benchmark models, tasks, and questions is publicly available, making modifications convenient (2) they include a wide variety of answer-generating models, tasks, and task questions for a comprehensive evaluation (3) their human evaluations validated the reliability of state-of-the-art judging models (GPT-4 and GPT-4-Turbo) on their evaluation instances, hence model untested by prior work, if reaching high agreement with these validated judges, can be perceived reliable as well.
G.1 Judges, Candidates, and Tasks
Judge
In this study, we choose seven GPT, four Claude, and three Gemini models as the judges. The specific versions for API call are specified as follows: o1-mini-2024-09-12 for o1-mini, gpt-4o-2024-05-13 for GPT-4o, gpt-4-1106-preview for GPT-4-Turbo, gpt-4-0613 for GPT-4, and gpt-3.5-turbo-1106 for GPT-3.5-turbo; claude-3-5-sonnet-20240620, claude-3-opus-20240229, claude-3-sonnet-20240229, and claude-3-haiku-20240307 for Claude series. The other model names and versions are as they are.
Model
The reference (or baseline) answer-generating models are vicuna-13b-v1.3 for MTBench and human for DevBench. They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations. The other models that are compared to the reference models, namely “Model" in our context, are listed as follows.
  • MTBench (30): alpaca-13b, baize-v2-13b, chatglm-6b, claude-instant-v1, claude-v1, dolly-v2-12b, falcon-40b-instruct, fastchat-t5-3b, gpt-3.5-turbo, gpt-4, gpt4all-13b-snoozy, guanaco-33b, guanaco-65b, h2ogpt-oasst-open-llama-13b, koala-13b, llama-13b, mpt-30b-chat, mpt-30b-instruct, mpt-7b-chat, nous-hermes-13b, oasst-sft-4-pythia-12b, oasst-sft-7-llama-30b, palm-2-chat-bison-001, rwkv-4-raven-14b, stablelm-tuned-alpha-7b, tulu-30b, vicuna-33b-v1.3, vicuna-7b-v1.3, wizardlm-13b, wizardlm-30b
  • DevBench (10): codellama-7b-instruct, codellama-13b-instruct, codellama-34b-instruct, deepseek-coder-1.3b-instruct, deepseek-coder-6.7b-instruct, deepseek-coder-33b-instruct, gpt-3.5-turbo-1106, gpt-4-0125-preview, gpt-4-0613, gpt-4-1106-preview
The model names are exactly what MTBench Zheng et al. (2024b) and DevBench Li et al. (2024a) used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
Task
For tasks, we also follow the original studies of these two benchmarks, except for DevBench we separate the gerenal metrics into detailed ones and considered them as different tasks. In this sense, our study experiments on the following tasks to provide a comprehensive study on the positon bias of LLM-as-a-Judge:
  • MTBench (8): coding, extraction, humanities, math, reasoning, roleplay, stem, and writing.
  • Devbench (14):
    • UML class (4): cohesion_and_decoupling, complexity, practicability, and faithfulness
    • UML sequence (5): cohesion_and_decoupling, interaction_complexity, practicability, uniformity_and_integration, and faithfulness
    • architecture design (5): conformance, design_and_coding, practicability, uniformity_and_integration, and faithfulness
G.2 Prompt Settings
We follow the original prompt settings of MTBench and DevBench in our study of pairwise comparative LLM-as-a-Judge.
Though written differently, these prompts all share same key components:
  • A system prompt explaining the judging task and the role the LLM should be playing.
  • Emphasized “should" and“shouldn’t"s.
  • A prompt structure with placeholders for specific questions and model answers
  • A specified output format for later judgment extraction
  • Chain-of-Thought Wei et al. (2022) prompts requiring the LLM judge to provide reasons for its judgment
The detailed prompt templates are specified below.